# Wandering around: A bioinspired approach to visual attention through object motion sensitivity

### v1

@@ 4.3 SNN Proto-object model: Need for proto-object saliency
Sub-dataset Event map [mitrokhin2019ev] OMS map Saliency map
Wall Refer to caption Refer to caption Refer to caption
Wall Refer to caption Refer to caption Refer to caption
Fast Refer to caption Refer to caption Refer to caption
Fast Refer to caption Refer to caption Refer to caption
Floor Refer to caption Refer to caption Refer to caption
Figure 11: From left to right: sub-dataset name, Event map, OMS map and saliency map for several representative examples from the EVIMO dataset [mitrokhin2019ev]. The colour legend for the saliency map follows the jet map, with blue and red indicating minimum and maximum values, respectively.
@@
Where $\mathbf{x}_{\text{max}}(t)$ represents the coordinates of the maximum value in the saliency map at time step $t$, $\mathbf{B}(t)$ is the bounding box created around $\mathbf{x}_{\text{max}}(t)$ (8$\times$8 pixels), $\mathbf{M}(t)$ is the object mask at time step $t$, and $\delta(\cdot)$ is an indicator function that returns 1 if the bounding box $\mathbf{B}(t)$ intersects with the mask $\mathbf{M}(t)$. Table 3 shows the mean accuracy percentages and standard deviations for each sub-dataset.
@@
Table 3: Accuracy of the SNN Proto-object model (in percent) for object detection is reported for each EVIMO sub-dataset [mitrokhin2019ev].
@@
To evaluate the system’s performance across a broader range of scenarios beyond the EVIMO dataset, which is confined to an office environment, the model has also been benchmarked against the Low-Light Event Video Object Segmentation (LLE-VOS) dataset [li2024event]. The LLE-VOS dataset comprises 70 video sequences, each containing both standard and low-light versions, along with segmentation labels and low-light event data. These sequences were recorded in diverse environments, including sports facilities, outdoor play areas, educational spaces, conference rooms, and wildlife enclosures, with some sequences involving camera motion and others without. The data are collected using two identical cameras with different exposure times to generate normal-light and low-light pairs of synthetically generated events.
@@
Sub-dataset Normal-light RGB [li2024event] Low-light RGB [li2024event] Event map [li2024event] Annotation [li2024event] OMS map Saliency map Accuracy %
00002 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 84.13
00011 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 89.88
00064 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 87.47
00033 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 72.96
00031 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 55.55
00025 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 47.84
Figure 12: From left to right: sub-dataset name, Normal-light RGB frame, Low-light RGB frame, Event map, Annotated frame, OMS map, saliency map and accuracy in percentages for several representative examples from the LLE-VOS dataset [li2024event]. The colour legend for the saliency map follows the jet map, with blue and red indicating minimum and maximum values, respectively.
Figure 12 shows the normal-light RGB frame, low-light RGB frame, event map, annotation, OMS map, saliency map, and accuracy percentages for examples from the LLE-VOS sub-datasets. The Figure presents six examples with varying degrees of accuracy to provide a better understanding of the system’s behaviour on the data. The system is able to segment motion in indoor and outdoor conditions (i.e. sub-dataset 11 and 33) reaching up to 89.88% accuracy generating an OMS map and saliency map coherent with the nature of the event-based data. Example 25 demonstrates the ability of the SNN Proto-object model to aggregate the response to moving objects while disregarding clutter (e.g., foliage). However, similar to Example 31, and coherently with the results on speed found in the characterisation of the system in Section 3.1.1, the camera motion in Example 25 is too rapid, causing the system to fail in accurately segmenting moving objects, resulting in accuracy rates of 47.84% and 55.55%, respectively. The final mean accuracy in detecting objects across all sub-datasets is 53.3%, with a standard deviation of 25.56, reflecting the large variability in the fast and slow camera motion sub-datasets. Although the system focuses on the accuracy detection of a single salient object, its performance is slightly lower than that reported in [li2024event] (67.8%), which employs a learning approach to object detection. This underscores the potential of the sOMS and SNN Proto-object model for future implementations.
Figure 13: Real-world demonstrator, the Bioinspired attention system. On the left, a view of the setup with the Speck device from Synsense mounted on a FLIR PTU, and the computer displaying the response view. On the right, the response view is shown, including the Events map, the sOMS map, sAC coordinates (where the red dot is the fovea, the green dot is the salient point $P(x,y)$ and the yellow dot sAC commands ( $u_{pan}$ $u_{tilt}$).
@@ 4.5 Real-world demonstration: Bioinspired attention system wandering around
The Bioinspired attention system (see Figure 13) was demonstrated in various environments to evaluate its robustness and performance. These environments included: (1) an indoor scenario with alternate hand-shaking to test stability, (2) an indoor scene with a person and animals moving dynamically, (3) an outdoor scene with challenging sunlight conditions, and (4) an indoor scene where a person was positioned at a distance from the sensor and waved at the camera. Across these scenarios, the system exhibited a mean latency of $0.124$ seconds to perform the object motion segmentation and object detection generating the salient point, demonstrating a rapid response to saccadic commands and bringing the desired salient point to the fovea. Specifically, $0.0042\pm 0.0036$ seconds (over 10 trials) for the object motion sensitivity module and the proto-object detection module, and $0.12\pm 0.06$ seconds for the visual attention module. The introduction of fixational eye movements, which are small adjustment movements compared to fast saccadic movements, is essential for performing object motion segmentation accurately and successfully detecting the next salient point of interest. The video showing the real-time demonstration can be found here: https://youtu.be/dcAJlDgVR0o.
@@ 5 Discussion
The characterisation of the sOMS model confirmed the sigma sizes proposed in [clerico2024retina]. However, the model exhibits a comparable mean MFR and mean ISI even with $\sigma_{c}=2$. Additionally, the mean MFR indicates increased activity for a larger kernel size ($s=32$), while maintaining a similar mean ISI. This suggests comparable performance and allows us to exclude the use of larger kernel sizes, as they result in higher MFR. The system performs qualitatively well across various scenarios with diverse spatial frequencies, confirming the model’s scale invariance. The characterisation also demonstrates that the model does not completely suppress activity in the Eye Only condition but effectively performs object motion segmentation, enhancing motion contrast in all Eye+Object cases where the foreground speed is greater than the background speed ($s_{F}>s_{B}$). From an active vision perspective, this highlights the necessity of small corrective movements, such as fixational eye movements, to effectively segment motion in the scene and detect the next point of interest, an aspect found to be essential for attentional shifts [gu2024microsaccades].
@@
The sOMS model effectively suppresses 85% of events while achieving up to a mean IoU of 82.24% and mean SSIM accuracy of 96% when compared to the ground truth, demonstrating that the majority of the suppressed events do not pertain to objects in motion. In addition, although the benchmark comparison of the mean IoU (%) indicates lower performance compared to [clerico2024retina], likely due to the sparse responses generated by the sCNN layer, our results demonstrate greater consistency across the entire dataset as demonstrated by the standard deviation of both mean IoU and mean SSIM. The model also demonstrates its effectiveness across diverse scenarios, as evidenced by the characterisation with grating stimuli and benchmarking across all the EVIMO sub-datasets. Akin to the event-based nature of the input, the object motion segmentation in the sOMS map generates a sparse mask that highlights the leading and trailing edges of objects moving in the scene. The sOMS model demonstrates consistent performance even with the LLE-VOS benchmark, which contains a variety of dynamic indoor and outdoor scenes. As expected from the sOMS characterisation, the model fails when the camera motion exceeds the speed of the dynamic entities in the scene. This further reinforces the idea that the model performs best with small sensor movement adjustments before detecting the salient object and performing a saccade toward the next one.
@@
The introduction of the SNN Proto-object model adds an additional processing layer to detect potential objects in the scene, further refining object motion segmentation and perceptually grouping the visual scene information. The SNN Proto-object implementation, adapted on the sinabs library from the previously proposed attention model [Dangelo_etal22], confirms its ability to accurately detect objects in the scene. The SNN Proto-object model, grounded in the principles of the Gestalt laws, facilitates the perceptual grouping of edge features with close contours, effectively disregarding lines and potential clutter. As shown in Figure 8, the SNN Proto-object model enhances the sOMS model’s ability to detect potential objects in the scene, achieving an accuracy up to 88.8% (as shown in Figure 11). Additionally, the sOMS generates the OMS map, producing thicker edge maps of moving objects and thereby facilitating proto-object detection. Consistent with the event-based nature of the sOMS input, the saliency map detects moving objects, reaching 89.88% accuracy in the LLE-VOS dataset, focusing attention on the most salient point in the scene. As shown by the high standard deviation across the entire dataset (25.56), integrating the sOMS model with the SNN Proto-object model is most effective in setups where movements consist of small adjustments to detect the maximum salient point and foveate toward it. From an active vision perspective, this aligns with the non-uniform distribution of the retinal structure, which concentrates high resolution in the center of the visual field—the fovea—to facilitate further object recognition.
@@
The sOMS model successfully performs multi-object motion segmentation and detection, achieving up to 96% accuracy, while the SNN Proto-object model achieves up to 88.8% accuracy in object detection and 89.9% accuracy under challenging lighting conditions. Both learning-free models demonstrate robust performance across a wide range of scenarios, effectively handling motion segmentation and detection without the need for computationally expensive training or adaptation to varying conditions. This opens the possibility of integrating learning algorithms for more complex tasks, such as object recognition.
@@
Thanks to the sAC control, the complete closed-loop real-time bioinspired attention system performs effectively, demonstrating robustness across various scenes. The additional fixational eye movements following the previous saccade (at the maximum salient point) facilitate precise adjustment movements, enabling reliable and consistent object motion segmentation in local motion. This suggests that the system detects objects in this context, primarily using saccades to transition from one salient point to the next. As indicated by Crevecoeur et al. [crevecoeur2017saccadic], the partial saccadic suppression may result from efficient sensorimotor computations. This aligns with our findings, demonstrating that the bioinspired attention system can accurately segment and detect object motion only when these small adjustment movements are incorporated, allowing saccades to be used for transitioning between salient points.
@@ 6 Conclusion
The entire bioinspired system demonstrates the capability to effectively perform object motion segmentation and saccade on the most salient proto-object. The characterisation of both models, the sOMS and the SNN Proto-object model provides a comprehensive evaluation of their benefits and limitations. The proposed complete end-to-end bioinspired attention system architecture effectively enhances motion contrast and prioritises coherent objects while disregarding background features, accurately performing saccades towards the salient point of the scene. The system, characterised using ideal gratings and benchmarked on EVIMO, achieves a maximum of 82.2% mean IoU and 96% mean SSIM in multi-object motion segmentation, with object detection accuracies up to 88.8% in office scenarios and 89.8% in low-light indoor/outdoor conditions, as evaluated on LLE-VOS. This architecture introduces an innovative visual attention system for object motion segmentation based solely on visual cues, maintaining high temporal resolution while eliminating the need for calibration and synchronisation with external sensors such as IMUs. This aligns with the lack of scientific evidence suggesting that retinal ganglion cells (RGCs) possess proprioceptive capabilities. Furthermore, the saliency-based attention model can perceptually group motion features, enabling the detection of potential salient objects coherently with the nature of the even-based input. The resulting attention mechanism, based on object motion segmentation, reveals a single salient point, requiring a shift of the gaze to place it in the fovea for further stable inspection with high resolution. Implementing the non-uniform structure of the retina, which is beyond the scope of this work, would enhance the consistency of the proposed system. As demonstrated by D’Angelo et al[d2020event], this would further reduce noise, decrease the number of retinal events, and enhance motion direction detection through the varying receptive field sizes from the periphery to the fovea, offering deeper insights into human selective attention. This highlights a crucial consideration regarding camera resolution: a high resolution is likely not necessary, as we can actively direct the gaze toward the point of interest. The central finding of this work is the introduction of fixational eye movements to effectively highlight object motion, enabling stable detection of the salient point while utilising the saccades solely for transitioning to the next salient point. Without fixational eye movements, the system would shift from one salient point to another without sufficient time to stabilise the visual field and detect a stable point of interest. The partial suppression of information during saccades, as demonstrated in our characterisation of the sOMS model, aligns with recent literature suggesting that information processing continues even during saccades, while perceptual stability is maintained. Specifically, it has been shown that saccades suppress visual processing of both color and luminance in the early visual cortex, but do not completely halt it [zhang2024execution]. Furthermore, presaccadic visual content can modulate postsaccadic processing, further indicating ongoing visual processing during eye movements [stankov2021during]. These findings suggest that our visual system partially maintains continuous processing during saccades while preserving perceptual stability and preventing disorientation from constant visual motion. This underscores the embodied nature of our system, where active perception, through the interplay of fixational eye movements and saccades, directly shapes visual processing, mirroring biological mechanisms that integrate motor actions with sensory input to enhance stability and perception. Central to future experiments will be the investigation of saccades and fixational eye speeds, exploring the limits of the system and the PTU to achieve smoother saccadic movements and even smooth pursuit. Moreover, even if the scene remains static, the system is able to wander around effectively, thanks to the fixational eye movements and attention mechanisms that generate a bottom-up saliency map of the scene. Additionally, small retinal shifts create depth-dependent motion patterns for objects at different distances, leading to spike variations that provide an initial cue for monocular depth perception. A further development could integrate recurrence between the border ownership layer and the grouping layer enhancing a more dense segmentation of the objects [d2024event].
@@
The entire architecture is built using spiking neural networks, representing the first closed-loop neuromorphic visual attention approach through object motion sensitivity. The decision to implement the visual and control components on separate neuromorphic frameworks is based on the fact that at the time of writing the entire pipeline could likely not be accommodated on a single platform. This approach allows for the distribution of the architectures across different hardware leveraging the neuromorphic intermediate representation [pedersen2024neuromorphic]. The full deployment of the architecture is beyond the scope of this work; however, this study serves as the first attempt to characterise end-to-end active visual attention and lays the foundation for its complete implementation on neuromorphic hardware to demonstrate the advantages of end-to-end neuromorphic systems in terms of reduced latency and power consumption. The architecture is entirely model-based and not dependent on training, allowing the system to handle diverse environments and scenarios, as demonstrated in various experiments, ranging from ideal gratings to office scenarios and challenging indoor and outdoor environments. However, learning could be employed to adapt the parameters to further optimise the system’s performance.

### v2

@@ 4.3 SNN Proto-object model: Need for proto-object saliency
Figure 11: From left to right: sub-dataset name, Event map, OMS map and saliency map for several representative examples from the EVIMO dataset [59]. The colour legend for the saliency map follows the jet map, with blue and red indicating minimum and maximum values, respectively.
@@
[…]
@@
Table 3: Accuracy of the SNN Proto-object model (in percent) for object detection is reported for each EVIMO sub-dataset [59].
@@
To evaluate the system’s performance across a broader range of scenarios beyond the EVIMO dataset, which is confined to an office environment, the model has also been benchmarked against the Low-Light Event Video Object Segmentation (LLE-VOS) dataset [60]. The LLE-VOS dataset comprises 70 video sequences, each containing both standard and low-light versions, along with segmentation labels and low-light event data. These sequences were recorded in diverse environments, including sports facilities, outdoor play areas, educational spaces, conference rooms, and wildlife enclosures, with some sequences involving camera motion and others without. The data are collected using two identical cameras with different exposure times to generate normal-light and low-light pairs of synthetically generated events. Notably, the events used in this benchmark are exclusively derived from low-light conditions, emphasizing the superior capability of event-driven cameras to perform effectively in such challenging environments, where traditional frame-based cameras often face significant limitations.
@@
Figure 12: From left to right: sub-dataset name, Normal-light RGB frame, Low-light RGB frame, Event map, Annotated frame, OMS map, saliency map and accuracy in percentages for several representative examples from the LLE-VOS dataset [60]. The colour legend for the saliency map follows the jet map, with blue and red indicating minimum and maximum values, respectively.
Figure 12 shows the normal-light RGB frame, low-light RGB frame, event map, annotation, OMS map, saliency map, and accuracy percentages for examples from the LLE-VOS sub-datasets. The Figure presents six examples with varying degrees of accuracy to provide a better understanding of the system’s behaviour on the data. The accuracy values are computed as described in Section 4.3. The system is able to segment motion in indoor and outdoor conditions (i.e. sub-dataset 11 and 33), reaching up to 89.88% accuracy, generating an OMS map and saliency map coherent with the nature of the event-based data. Example 25 demonstrates the ability of the SNN Proto-object model to aggregate the response to moving objects while disregarding clutter (e.g., foliage). However, similar to Example 31, and coherently with the results on speed found in the characterisation of the system in Section 3.1.1, the camera motion in Example 25 is too rapid, causing the system to fail in accurately segmenting moving objects, resulting in accuracy rates of 47.84% and 55.55%, respectively. The final mean accuracy in detecting objects across all sub-datasets is 53.3%, with a standard deviation of 25.56, reflecting the large variability in the fast and slow camera motion sub-datasets. Although the system focuses on the accuracy detection of a single salient object, its performance is slightly lower than that reported in [60] (67.8%), which employs a learning approach to object detection. This underscores the potential of the sOMS and SNN Proto-object model for future implementations.
Figure 13: Real-world demonstrator, the Bioinspired attention system. On the left, a view of the setup with the Speck device from Synsense mounted on a FLIR PTU, and the computer displaying the response view. On the right, the response view is shown, including the Events map, the sOMS map, sAC coordinates (where the red dot is the fovea, the green dot is the salient point $P(x,y)$ and the yellow dot sAC commands ( $u_{pan}$ & $u_{tilt}$).
Figure 14: From left to right: time window ($t_{w}$) and Saliency map time window ($St_{w}$, Event map, OMS map and Saliency map at $t_{w}$. Black dots on the saliency map represent the salient points, and the yellow lines the saccades trajectories.
@@ 4.5 Real-world demonstration: Bioinspired attention system wandering around
The Bioinspired attention system (see Figure 13) was demonstrated in various environments to evaluate its robustness and performance. These environments included: (1) an indoor scenario with alternate hand-shaking to test stability, (2) an indoor scene with a person and animals moving dynamically, (3) an outdoor scene with challenging sunlight conditions, and (4) an indoor scene where a person was positioned at a distance from the sensor and waved at the camera. Across these scenarios, the system exhibited a mean latency of $0.124$ seconds to perform the object motion segmentation and object detection, generating the salient point, demonstrating a rapid response to saccadic commands and bringing the desired salient point to the fovea. Specifically, $0.0042\pm 0.0036$ seconds (over 10 trials) for the object motion sensitivity module and the proto-object detection module, and $0.12\pm 0.06$ seconds for the visual attention module. The introduction of fixational eye movements, which are small adjustment movements compared to fast saccadic movements, is essential for performing object motion segmentation accurately and successfully detecting the next salient point of interest. The video showing the real-time demonstration can be found here: https://youtu.be/dcAJlDgVR0o.
4.5.1 Bioinspired attention system wandering around: evaluating the scanpath
To further validate the proposed architecture in robotic scenarios and underscore the necessity of spatiotemporal integration, we evaluated the system in a static setting involving objects placed on a shelf. Events were generated using the ATIS cameras mounted on the iCub robot, which executed eye movements through the motorised actuation of the “eye muscles,” thereby mimicking fixational eye movements. The eye movements covered 72 discrete positions along a circular trajectory, with updates occurring every 5 ms, resulting in a total duration of 360 ms and a pan-tilt angle of $1^{\circ}$ per position. During this process, we recorded the system’s output, including the maximum salient points, saliency maps, and simulated saccades (representing the scanpath, namely the collection of eye trajectories). We considered two distinct operational conditions. In the first case, Saccades, we assumed no time for spatiotemporal integration. The system computed the Event map, OMS map, and Saliency map at every time window $t_{w}=10$ ms, simulated the saccade, and subsequently generated the final cumulative saliency map using a saliency map time window $St_{w}$. In the second case, Fixational eye movements, we set $t_{w}=St_{w}$ = 124 ms (same latency of the real-world demonstrator), allowing the system to integrate information temporally within a single, consistent window. The responses in Figure 14 further underscore the need for spatiotemporal integration to enable meaningful object detection (Fixational eye movements case) and prevent random saccadic jumps to non-informative locations (Saccades case), while also demonstrating the model’s effectiveness in static scenarios. This analysis lays the foundation for future research on robotic eye movements for event-based active vision.
@@ 4.5.1 Bioinspired attention system wandering around: evaluating the scanpath
The code to reproduce the data for the characterisation is available at https://github.com/GiuliaDAngelo/Speckegomotion/blob/main/CreateStimuli.py.
All the data used in the study can be easily reproduced by running the scripts provided in our GitHub repository. Furthermore, the two benchmark datasets employed are already publicly accessible online (EVIMO [59] & LLE-VOS [60]).
@@ 5 Discussion
The characterisation of the sOMS model confirmed the sigma sizes proposed in [52]. However, the model exhibits a comparable mean MFR and mean ISI even with $\sigma_{c}=2$. Additionally, the mean MFR indicates increased activity for a larger kernel size ($s=32$), while maintaining a similar mean ISI. This suggests comparable performance and allows us to exclude the use of larger kernel sizes, as they result in higher MFR. The system performs qualitatively well across various scenarios with diverse spatial frequencies, confirming the model’s scale invariance. The characterisation also demonstrates that the model does not completely suppress activity in the Eye Only condition but effectively performs object motion segmentation, enhancing motion contrast in all Eye+Object cases where the foreground speed is greater than the background speed ($s_{F}>s_{B}$).
@@
The sOMS model effectively suppresses 85% of events while achieving a mean IoU of up to 82.24% and a mean SSIM accuracy of 96% in multi-object motion segmentation compared to the ground truth, demonstrating that the majority of the suppressed events are not related to objects in motion. In addition, although the benchmark comparison of the mean IoU (%) indicates lower performance compared to [52], likely due to the sparse responses generated by the sCNN layer, our results demonstrate greater consistency across the entire dataset as demonstrated by the standard deviation of both mean IoU and mean SSIM. The model also demonstrates its effectiveness across diverse scenarios, as evidenced by the characterisation with grating stimuli and benchmarking across all the EVIMO sub-datasets. Akin to the event-based nature of the input, the object motion segmentation in the sOMS map generates a sparse mask that highlights the leading and trailing edges of objects moving in the scene. The sOMS model demonstrates consistent performance even with the LLE-VOS benchmark, which contains a variety of dynamic indoor and outdoor scenes. As expected from the sOMS characterisation, the model fails when the camera motion exceeds the speed of the dynamic entities in the scene. This further reinforces the idea that the model performs best with small sensor movement adjustments before detecting the salient object and performing a saccade toward the next one.
@@
The introduction of the SNN Proto-object model adds an additional processing layer to detect potential objects in the scene, further refining object motion segmentation and perceptually grouping the visual scene information. The SNN Proto-object implementation, adapted on the sinabs library from the previously proposed attention model [39], confirms its ability to accurately detect objects in the scene. The SNN Proto-object model, grounded in the principles of the Gestalt laws, facilitates the perceptual grouping of edge features with close contours, effectively disregarding lines and potential clutter. As shown in Figure 8, the SNN Proto-object model enhances the sOMS model’s ability to detect potential objects in the scene, achieving an accuracy up to 88.8% (as shown in Figure 11). Additionally, the sOMS generates the OMS map, producing thicker edge maps of moving objects and thereby facilitating proto-object detection. Consistent with the event-based nature of the sOMS input, the saliency map detects moving objects with 89.88% accuracy on the LLE-VOS dataset under challenging lighting conditions, effectively focusing attention on the most salient point in the scene. As shown by the high standard deviation across the entire dataset (25.56), integrating the sOMS model with the SNN Proto-object model is most effective in setups where movements consist of small adjustments to detect the maximum salient point and foveate toward it. From an active vision perspective, this aligns with the non-uniform distribution of the retinal structure, which concentrates high resolution in the center of the visual field—the fovea—to facilitate further object recognition.
@@
Both learning-free models demonstrate robust performance across a wide range of scenarios, effectively handling motion segmentation and detection without the need for computationally expensive training or adaptation to varying conditions. This opens the possibility of integrating learning algorithms for more complex tasks, such as object recognition.
@@
Thanks to the sAC control, the complete closed-loop real-time bioinspired attention system performs effectively, demonstrating robustness across various scenes. The additional fixational eye movements following the previous saccade (at the maximum salient point) facilitate precise adjustment movements, enabling reliable and consistent object motion segmentation in local motion. This suggests that the system detects objects in this context, primarily using saccades to transition from one salient point to the next. As indicated by Crevecoeur et al. [90], the partial saccadic suppression may result from efficient sensorimotor computations. This aligns with our findings, demonstrating that the bioinspired attention system can accurately segment and detect object motion only when these small adjustment movements are incorporated, allowing saccades to be used for transitioning between salient points.
@@ 6 Conclusion
The proposed end-to-end bioinspired attention system demonstrates robust performance in segmenting object motion during fixational eye movements and saccades toward the most salient proto-object. Detailed characterisation and benchmarking of the sOMS and SNN Proto-object models show that the system enhances motion contrast, prioritises coherent objects, and disregards background features across varied scenarios, including office environments and low-light indoor/outdoor scenes.
This architecture introduces an innovative visual attention system through object motion segmentation that relies exclusively on visual cues, maintaining high temporal resolution without requiring calibration or synchronization with external sensors such as IMUs. This design aligns with the absence of proprioceptive capabilities in retinal ganglion cells (RGCs), as there is no scientific evidence to suggest that RGCs possess proprioceptive functions. Consequently, proprioceptive feedback from external devices is not necessary for effective object motion segmentation during fixational eye movements. Furthermore, the saliency-based attention model can perceptually group motion features, enabling the detection of potential salient objects coherently with the nature of the even-based input. The resulting attention mechanism, based on object motion segmentation, reveals a single salient point, requiring a shift of the gaze to place it in the fovea for further stable inspection with high resolution. Implementing the non-uniform structure of the retina, which is beyond the scope of this work, would enhance the consistency of the proposed system. As demonstrated by D’Angelo et al[40], this would further reduce noise, decrease the number of retinal events, and enhance motion direction detection through the varying receptive field sizes from the periphery to the fovea, offering deeper insights into human selective attention. This highlights a crucial consideration regarding camera resolution: a high resolution is likely, not necessary, as we can actively direct the gaze toward the point of interest.
From an active vision perspective, the central finding of this work is the need for small corrective movements, such as fixational eye movements, for spatiotemporal integration. These movements effectively highlight object motion, enabling the stable detection of salient points, which is crucial for attentional shifts [91], while reserving saccades solely for transitioning to the next salient point. Without fixational eye movements, the system would shift abruptly from one salient point to another, failing to stabilise the visual field and hindering the detection of stable points of interest. The partial suppression of information during saccades, as demonstrated in our characterisation of the sOMS model, aligns with recent literature suggesting that information processing continues even during saccades, while perceptual stability is maintained. Specifically, it has been shown that saccades suppress visual processing of both color and luminance in the early visual cortex, but do not completely halt it [92]. Furthermore, presaccadic visual content can modulate postsaccadic processing, further indicating ongoing visual processing during eye movements [93]. These findings suggest that our visual system partially maintains continuous processing during saccades while preserving perceptual stability and preventing disorientation from constant visual motion.
This underscores the embodied nature of our system, where active perception, through the interplay of fixational eye movements and saccades, directly shapes visual processing, mirroring biological mechanisms that integrate motor actions with sensory input to enhance stability and perception. Central to future experiments will be the investigation of saccades and fixational eye speeds, exploring the limits of the system and the PTU to achieve smoother saccadic movements and even smooth pursuit. Moreover, even if the scene remains static, the system is able to wander around effectively, thanks to the fixational eye movements and attention mechanisms that generate a bottom-up saliency map of the scene. Additionally, small retinal shifts create depth-dependent motion patterns for objects at different distances, leading to spike variations that provide an initial cue for monocular depth perception. A further development could integrate recurrence between the border ownership layer and the grouping layer enhancing a more dense segmentation of the objects [94].
@@
The entire architecture is built using spiking neural networks, representing the first closed-loop neuromorphic visual attention approach through object motion sensitivity. The decision to implement the visual and control components on separate neuromorphic frameworks is based on the fact that at the time of writing the entire pipeline could likely not be accommodated on a single platform. This approach allows for the distribution of the architectures across different hardware leveraging the neuromorphic intermediate representation [95]. The full deployment of the architecture is beyond the scope of this work; however, this study serves as the first attempt to characterise end-to-end active visual attention and lays the foundation for its complete implementation on neuromorphic hardware to demonstrate the advantages of end-to-end neuromorphic systems in terms of reduced latency and power consumption. The architecture is entirely model-based and not dependent on training, allowing the system to handle diverse environments and scenarios, as demonstrated in various experiments, ranging from ideal gratings to office scenarios and challenging indoor and outdoor environments. However, learning could be employed to adapt the parameters to further optimise the system’s performance.